OpenStudy (anonymous):

how to get a perimeter a of a rectangle with a width of 8x^2y cm and 12x^2

OpenStudy (anonymous):

perimeter=2(8x^2y)+2(12x^2)=16x^2y+24x^2 with the information given this is the most you can do

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

uh

OpenStudy (anonymous):

:})

OpenStudy (anonymous):

are u sure??

OpenStudy (anonymous):

my teacher wrote 40x^2cm

OpenStudy (anonymous):

in your first term you have a y variable (look at your orignial question); if that is not here then you get the answer 40x^2

OpenStudy (anonymous):

perimeter=2(8x^2)+2(12x^2)=16x^2+24x^2=40x^2

OpenStudy (anonymous):

with out the other y 40x^ is the answer??

OpenStudy (anonymous):

if the y is there the two terms are not "similar" or "like" and we can't add them together and make a single term

OpenStudy (anonymous):

ahh kk thnx

OpenStudy (anonymous):

np :})

OpenStudy (anonymous):

pls dont leave i have more questions ... sori

OpenStudy (anonymous):

uh how to do this x^3.x^2=x^6 the solution pls.......